Three genes with recessive mutant phenotypes are located on the same chromosome in maize: waxy endosperm (wx), shrunken endosperm (sh), and virescent seedling (v). A plant with all three mutant phenotypes is crossed with a true-breeding, wild-type plant. Resulting F1 plants are then backcrossed to the parental mutant line to produce the following progeny:

Waxy Shrunken Virescent Number
wx sh + 89
+ + v 92
+ + + 1482
wx sh v 1475
+ sh + 480
wx + v 477
wx + + 290
+ sh v 282

Draw a chromosome map with the correct distances (rounded to the nearest cM) noted between

adjacent pairs of genes.  

Chromosome map (horizontal):
Answer #1

Here, wx wx sh sh v v(recessive mutant for all 3)..............X..............................wx+wx+ sh+sh+ v+v+(wild type)

Gametes ....wx sh v....................................................................................................wx+ sh+ v+

F1 ....................................wx+wx sh+sh v+v

Now the F1 progeny is crossed with recessive parent (test-cross)

Hence, wx+wx sh+sh v+v..........................X........................................................................................wx wx sh sh v v

Gametes: wx+ sh+v+,wx+shv,wx+sh+v,wx+shv+,wx sh v, wx sh+v+, wx sh+v, wx sh v+.....................wx sh v

F1 : wx+wx sh+sh v+v =(+++)= 1482

wx wx sh sh v v = (wx sh v) = 1475

wx+ wx sh sh v v= ( + sh v) = 282

wx+ wx sh+ sh v v= ( + + v) = 92

wx+ wx sh sh v+v = (+ sh +) = 480

wx wx sh+ sh v+ v= (wx + +)= 290

wx wx sh+ sh v v= ( wx + v) = 477

wx wx sh sh v+v = ( wx sh +)= 89

·         Largest number of offsprings: parental genotype of the heterozygous F1 plant =

wx+wx sh+sh v+v =(+++)= 1482 & wx wx sh sh v v = (wx sh v) = 1475

·         Smallest number of offsprings: double cross-over leading to these genotypes =

wx wx sh sh v+v = ( wx sh +)= 89 & wx+ wx sh+ sh v v= ( + + v) = 92

ORDER of gene: wx.v.sh (double crossin gover shown in diagram)

wx+ v+ sh+ = 1482

wx v sh = 1475

wx+ v sh = 282

wx v+ sh+ = 290

wx+ v+ sh = 480

wx v sh+ = 477

wx+ v sh+ = 92

wx v+ sh = 89

Now, Distance between wx & v = ( Single crossing over between wx & v + Double crossing over )/ Total =

((282+290)+ (89+92))/ 4667 = 0.161 or 16.1 cM

Now, Distance between v & sh = ( Single crossing over between v & sh + Double crossing over )/ Total =

((477+480)+(89+92))/4667 = 0.244 or 24.4 cM

Distance between wx & sh = ( Single crossing over between wx & sh + 2 X Double crossing over )/ Total =

((282+290+477+480)+ 2 X (89+92))/4667 = 0.405 or 40.5cM

Hence the chromosomal map is :